You were asked to find the p-value associated with the test statistic, given the following information:

A medical researcher wants to compare the pulse rates of smokers and non-smokers. He believes that the pulse rate for smokers and non-smokers is different and wants to test this claim at the 0.01 level of significance. A sample of 74 smokers has a mean pulse rate of 87, and a sample of 76 non-smokers has a mean pulse rate of 8484. The population standard deviation of the pulse rates is known to be 88 for smokers and 77 for non-smokers.

In the previous step, you determined that to find the p-value, you need the area under the normal curve to the left of z= −2.44 and to the right of z= 2.44. Using the normal distribution table or technology, enter the p-value for this hypothesis test below.
Answer #1

p-value = P(Z < -2.44) + P(Z > 2.44) = 0.0073 + 0.0073 = 0.0146